if 5% of the energy supplied to an incandescent light bulb is radiated as visible light, how many visible light photons are emitted by 100 watt bulb? assume wavelength of all visible photons to be 5600*10 raise to the power -10.
 

Grade:12

1 Answers

Arun
6 years ago
Dear Kashish
 
\lambda = 5600 A
Energy of one photon is E = h\nu = h*c/\lambda
E = 6.63* 10^-34 * 3* 10^8/(5600* 10^-10) = 3.616 * 10^-19 J
A 100 watt bulb requires 100 joule of energy per second.
 
Energy radiated by the bulb as visible light per second = 5% of 100
= 5 Joule
 
Number of photons emitted per second=
Energy radiated per sec/ energy of one ohoton
= 5/(3.616 * 10^ -19) 
= 1.38 * 10^19
 
Regards
Arun (askIITians form expert)